how do you type the less than/equal to sign on the computer? anyway, here's the question. Hope you can understand:
solve for x: 1/5x - 2 (is greater than or equal to) 3/4x +5

Answer by stanbon(75887)

1/5x - 2 >= 3/4x +5
(1/5)x-(3/4)x = 5+2
(4/20)x - (15/20)x = 7
(-11/20)x=7
-11x=140
x=-140/11
Cheers,
Stan H.